Solve
p^2 -2xyp+4y^2=0,where p=dy/dx
1
Expert's answer
2020-10-14T17:03:21-0400

Given,

"p^2 -2xyp+4y^2=0"

Now, solve the quadratic equation in p, thus we get,


"p=\frac{2xy\pm \sqrt{4x^2y^2-16y^2}}{2}\\
\implies p=y(x\pm\sqrt{x^2-4})"

Hence,


"\frac{dy}{dx}=y(x\pm\sqrt{x^2-4})\\
\implies 
\frac{dy}{y}=(x\pm\sqrt{x^2-4})dx\\
\implies \int \frac{dy}{y}=\int(x\pm\sqrt{x^2-4})dx\\
\implies \ln y=\int xdx\pm\int \sqrt{x^2-4})dx\\
\implies \ln y=\frac{x^2}{2}\pm \int\sqrt{x^2-4})dx"

Since,

we know from standard integral that


Thus,


"\int \sqrt{x^2-4}dx=\dfrac{x\sqrt{x^2-4}}{2}-2\ln\left(\left|\sqrt{x^2-4}+x\right|\right)+C"

Hence,


"\ln y=\frac{x^2}{2}\pm \dfrac{x\sqrt{x^2-4}}{2}-2\ln\left(\left|\sqrt{x^2-4}+x\right|\right)+C\\
\implies y=e^{\frac{x^2}{2}\pm \dfrac{x\sqrt{x^2-4}}{2}-2\ln\left(\left|\sqrt{x^2-4}+x\right|\right)+C}\\
\implies y=C_1e^{\frac{x^2}{2}\pm \dfrac{x\sqrt{x^2-4}}{2}-2\ln\left(\left|\sqrt{x^2-4}+x\right|\right)}"